Onslow Co. purchases a used machine for $288,000 cash on January 2 and readies it for use the next day at a $8,000 cost. On January 3, it is installed on a required operating platform costing $1,600, and it is further readied for operations. The company predicts the machine will be used for six years and have a $34,560 salvage value. Depreciation is to be charged on a straight-line basis. On December 31, at the end of its fifth year in operations, it is disposed of. 2. Prepare journal entries to record depreciation of the machine at December 31
Answer #1

Answer

  • Working

A = 288000 + 8000 + 1600

Cost

$          297,600.00

B

Residual Value

$            34,560.00

C=A - B

Depreciable base

$          263,040.00

D

Life [in years]

6

E=C/D

Annual SLM depreciation

$            43,840.00

  • Requirement 2

Date

Accounts title

Debit

Credit

31 Dec

Depreciation expense - Equipment

$      43,840.00

Accumulated Depreciation - Equipment

$ 43,840.00

(depreciation expense recorded)